OpenStudy (anonymous):

c^2-10c=0 what rae the solutions of the equation?

Parth (parthkohli):

\( \color{Black}{\Rightarrow c(c - 10) = 0}\) Use the zero-product rule :)

Parth (parthkohli):

So, the roots can be obtained by solving the equations: \( \color{Black}{\Rightarrow c = 0}\) \( \color{Black}{\Rightarrow c - 10 = 0}\)

OpenStudy (anonymous):

take c common c(c-10)=0 so c=0 c=-10=0 c=10 c=0 and c=10 are solutions